### v1

# Can stellar physics explain GW231123?

Abstract
The gravitational wave event GW231123 detected by the LIGO interferometers during their fourth observing run features two black holes with source-frame masses of $137^{+22}_{-17}{\rm M}_{\odot}$ and $103^{+20}_{-52}{\rm M}_{\odot}$ — well within or above the pair-instability black hole mass gap predicted by standard stellar evolution theory. Both black holes are also inferred to be rapidly spinning ($\chi_{1}\simeq 0.9$, $\chi_{2}\simeq 0.8$). The primary object in GW231123 is the heaviest stellar mass black hole detected to date, which, together with its extreme rotation, raises questions about its astrophysical origin. Accounting for the unusually large spin of $\sim 0.9$ with hierarchical mergers requires some degree of fine tuning. We investigate whether such a massive, highly spinning object could plausibly form from the collapse of a single rotating massive star. We simulate stars with an initial core mass of $160\,{\rm M}_{\odot}$ — sufficient to produce BH masses at the upper edge of the 90% credible interval for $m_{1}$ in GW231123 — across a range of rotation rates and $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rates. We find that: (i) rotation shifts the pair-instability mass gap to higher masses, introducing a significant ingredient that correlates masses and spins in gravitational wave predictions; and (ii) highly spinning BHs with masses $\gtrsim 150\rm M_{\odot}$ can form above the mass gap, implying that stellar evolution alone is sufficient to explain GW231123. Our results suggest that the primary object of GW231123 may be the first directly observed black hole that formed via direct core collapse following the photodisintegration instability.
preprint: IPPP/25/52
Introduction
On 23 November 2023, the LIGO Hanford and Livingston interferometers detected a short, $\sim 5$‐cycle signal consistent with the merger of two black holes (BHs) with source‐frame masses
$m_{1}=137^{+22}_{-17}\,{\rm M}_{\odot},\quad m_{2}=103^{+20}_{-52}\,{\rm M}_{\odot},$ (1)
[…]
Stellar evolution theory predicts a “pair‐instability” black hole mass gap (BHMG) between approximately $60$ and $130\,{\rm M}_{\odot}$, driven by electron–positron pair production in helium cores that powers (pulsational) pair‐instability supernovae (hereafter (P)PISN) and completely disrupts stars with core masses up to $\sim 135{\rm M}_{\odot}$ [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Various stellar parameters and other effects determine the location of the BHMG, most importantly the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate [15, 21, 16, 17, 18], which can give variations of $\sim 47{\rm M}_{\odot}$ [17] in the location of the upper and lower edges of the BHMG.
If it had zero spin, the primary BH in GW231123 would likely lie above the BHMG unless the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate were $\sim 3\sigma$ smaller than its median value [17]. However, this is based on simulations of (P)PISN which have neglected rotation, so this conclusion is unlikely to apply to GW231123 and its extremely high spins. Indeed, previous theoretical work [22] has shown that rotation can raise the lower edge of the gap by $\sim 15\%$ (from $45.5{\rm M}_{\odot}$ to $52.4{\rm M}_{\odot}$). It is currently not known how rotation affects the upper edge of the BHMG. These considerations raise the question of whether stellar evolution can produce a BH as massive and highly spinning as the primary in GW231123.
Core collapse instabilities
The predicted BHMG is a direct result of the physics of pair‐instability in massive helium cores. At core temperatures $\gtrsim 10^{9}\,\mathrm{K}$, high‐energy photons convert into non-relativistic electron–positron pairs, stealing pressure support and causing the core to contract. In turn, this contraction ignites explosive nuclear burning, primarily of oxygen. If the released nuclear energy exceeds the gravitational binding energy, the star is completely disrupted with no BH remnant. In a slightly lower mass range the explosion is weaker and instead ejects the outer layers in pulses (a PPISN), removing mass until the remaining core collapses to a lighter BH. Above a critical helium core mass ($\sim 130{\rm M}_{\odot}$), stellar cores become so hot ($T_{c}\gtrsim 9\times 10^{9}\,\mathrm{K}$) that high‐energy photons begin to photodisintegrate iron‐group nuclei into $\alpha$‐particles and free nucleons. This endothermic process robs the core of pressure support, accelerating collapse — a phenomenon known as the photodisintegration instability. In this scenario, the star implodes directly to a BH, defining the upper edge of the BHMG.
As noted above, the effects of rotation on the location of the upper edge of the BHMG are unknown. In general, the extra support provided by the centrifugal force results in lower central densities and temperatures compared with non-rotating stars [22, 23], so it can be expected that rotation delays the onset of photodisintegration instability and shifts the upper edge of the BHMG to heavier masses.
[…]
As the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate is raised, a larger fraction of 12C is converted into 16O during core helium burning. This higher oxygen abundance results in a stronger explosion, whilst the lower 12C abundance implies less carbon is available to form a convective 12C burning shell to counteract contraction [16]. Consequently, both the lower and upper edges of the mass gap move downward in tandem, preserving a nearly constant gap width of $\Delta M_{\rm BHMG}\approx 80^{+9}_{-5}{\rm M}_{\odot}$ [17]. Across the full $\pm 3\sigma$ uncertainty in the reaction rate, the lower boundary shifts from $\sim 59^{+34}_{-13}{\rm M}_{\odot}$ down to lighter values, while the upper boundary moves from $\sim 139^{+30}_{-14}{\rm M}_{\odot}$ to correspondingly lower masses [17].
Simulation suite
To assess the impact of rotation and nuclear physics on the boundaries of the pair-instability regime, we compute a grid of stellar models with initial helium core mass $160{\rm M}_{\odot}$. This value was chosen such that, if no PISN occurs, the star can be expected to form a BH with mass at the edge of the measured $90\%$ confidence for $m_{1}$ in GW231123. We vary the stellar rotation rate $\Omega/\Omega_{\mathrm{crit}}$ between 0 and $1$ to capture the effects of increasing centrifugal support. The critical angular velocity
$\Omega_{\rm crit}\equiv\sqrt{\frac{(1-\Gamma)\,G\,M}{R_{\rm eq}^{3}}}\,$, (2)
is defined as the rotation rate at which the outward centrifugal acceleration at the stellar equator exactly balances the inward effective gravitational acceleration, where $M$ is the gravitational mass, $R_{\rm eq}$ the equatorial radius, and $\Gamma\equiv L/L_{\rm Edd}=L\kappa/{4\pi GMc}$ is the local Eddington factor which accounts for radiation pressure. Additionally, we explore the variation of the rate of the temperature-dependent $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction over $\pm 3\sigma$ from its median rate $R_{\rm med}(T)$. Here, $\sigma$ parametrizes the variation as $R_{\sigma}(T)=R_{\rm med}(T)\exp[\sigma\mu(T)]$ with $\mu(T)$ the uncertainty at each temperature point, which is assumed to follow a log-normal distribution [24, 16, 17].
Our simulations were performed using the stellar structure code MESA, version 15140 [29, 30, 31, 32, 33, 34]. While MESA is a one-dimensional code — meaning that it assumes spherical symmetry — it is capable of simulating rotating objects via the shellular approximation, where the radial coordinate $r$ is replaced with the volume-equivalent radius of an isobar. A comprehensive description of rotation in MESA is provided in Ref. [33]. The technical details of our simulations are given in the End Matter, and our code will be made public upon publication of this work. We do not include the Spruit-Tayler (ST) dynamo [35], which allows for efficient differential rotation and retention of angular momentum throughout the evolution. The ST dynamo is commonly invoked in stellar models to enforce near-rigid rotation by coupling angular momentum between stellar layers [36, 37]. However, its physical validity remains debated, with some studies raising doubts about both the efficiency and even the onset of the magnetic instabilities required to close the dynamo loop under realistic stellar conditions (e.g., [38, 39]). The high spins inferred for both components of GW231123 are difficult to reconcile with models that include the ST dynamo, which predict nearly non-rotating BHs [22]; this suggests that angular momentum transport in massive stars may be much less efficient than the ST prescription assumes.
Figure 1: Final BH mass as a function of initial rotation with initial helium core mass $160\rm{\rm M}_{\odot}$, for several values of the $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate. The top panel shows results as a function of the initial stellar rotation $\Omega$ in units of the critical value $\Omega_{\rm crit}$; the bottom panel shows results as a function of the BH spin $\chi$. Stars with faster rotation and smaller $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ reaction rate undergo PISN and are not shown in the figure. The shaded contours indicate the $90\%$ credible intervals of the primary mass (orange) and spin (purple) in GW231123 [1]
Some simulations feature spin parameters exceeding unity at core-collapse. Although the BH may form with maximal spin, further accretion can reduce its spin if the added mass is not accompanied by sufficient angular momentum, or if excess angular momentum is expelled by feedback to avoid violating the Kerr limit. To estimate the final spin of the BH, we adopt a version of the model developed in Ref. [40], consistent with previous work [22]. We assume that the innermost $3{\rm M}_{\odot}$ of the progenitor collapses promptly into a BH with maximal spin. We have explicitly verified that the results are insensitive to the size of this core, see End Matter for more details. The remainder of the star is assumed to either fall directly into the BH or accrete through a disk, delivering additional mass and angular momentum. Accretion is regulated by feedback, ensuring that the final spin does not exceed the Kerr limit ($\chi\leq 1$).
Impact of rotation on the black hole mass gap
We show the results of our simulations in Fig. 1. The top panel shows the final BH mass as a function of the initial rotation rate $\Omega/\Omega_{\rm crit}$ for a $160\,M_{\odot}$ helium core, with $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rates varied from $2\sigma$ to $3\sigma$. The $3\sigma$ simulations produced PISN with no BH remnants for all $\Omega/\Omega_{\rm crit}$ (including zero) and are thus not shown. At low $\Omega/\Omega_{\rm crit}$ all models undergo direct collapse, yielding $M_{\rm BH}\approx 159\,M_{\odot}$. At some critical rotation threshold $\Omega_{\rm PISN}$, the pair‐instability disrupts the star completely and no remnant is left. This threshold increases from $\Omega_{\rm PISN}\sim 0.1\Omega_{\rm crit}$ for the $2\sigma$ rate to $\Omega_{\rm PISN}\sim 0.5\Omega_{\rm crit}$ for the $3\sigma$ rate, demonstrating that stronger carbon burning stabilizes the core against pair‐instability to higher rotation.
The bottom panel of Fig. 1 presents these same models in the BH spin–mass plane. Stars with a helium core mass of $160{\rm M}_{\odot}$ cannot form BHs with a spin above a certain threshold, which depends on the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate. In particular, BHs with spins and masses compatible with the $90\%$ confidence interval of the primary in GW231123 cannot be formed if the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate is smaller than its measured median value.
As anticipated above, our results can be explained by a delay in the onset of photodisintegration instability for fast rotation. As shown in Fig. 2, rotating stars reach lower core temperatures $T_{c}$ before the onset of explosive oxygen burning leads to rapid expansion and a temperature drop. Thus, these stars avoid reaching the conditions required for photodisintegration-induced collapse, and instead undergo PISN. For the median $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate, we find that this occurs at $\Omega_{\rm PISN}\sim 0.4\Omega_{\rm crit}$.
Figure 2: Central temperature $T_{c}$ versus central density $\rho_{c}$ for the simulations described in text, with $M_{\rm in}=160{\rm M}_{\odot}$ for the median $^{12}\mathrm{C}(\alpha,\gamma)^{16}\mathrm{O}$ rate. Non-rotating models reach higher core densities and temperatures, crossing into the regime where photodisintegration reactions lead to gravitational collapse, here assumed to be $T_{c}=9\times 10^{9}\rm\,K$. Rotating models, by contrast, terminate at lower $T_{c}$ and $\rho_{c}$, thereby avoiding collapse and instead undergoing a PISN.
Implications for GW231123
Our simulations imply that it is plausible to interpret GW231123’s primary component ($m_{1}\simeq 137{\rm M}_{\odot}$) as having formed via photodisintegration instability collapse. A similar claim was previously made [41] for the primary BH in GW190521, the former high‐mass record holder among GW events, although that interpretation depends strongly on the adopted prior distribution when analyzing the underlying GW data.
In our models, the $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ reaction rate needed to form GW231123’s primary is both broad and realistic, ranging from the median to rates $3\sigma$ stronger. The upper end is compatible with previous hints that higher rates are needed to explain GW observations e.g., the $35{\rm M}_{\odot}$ peak in the BH mass function [28].
Ostensibly, the combination of large masses and large spins in GW231123 could be explained by binaries formed through hierarchical mergers (for context, see Ref. [42] and references therein) and a few studies have already explored this scenario [1, 2, 4]. We argue that it is somewhat unlikely. Hierarchical mergers in dense stellar environments produce a spin distribution that is sharply peaked at $\chi\sim 0.7$ [43, 44], which matches the BH spin in GW231123 only near the lower edge of its 90% credible interval. Spins of $\chi_{1}\sim 0.9$ are too large to be explained as merger remnants [45], requiring substantial fine‐tuning of the progenitor binary (involving either a somewhat extreme mass ratio and/or mostly aligned spins). Our analysis shows that stellar evolution alone can naturally explain the primary BH in GW231123 without requiring a hierarchical-merger origin.
Outlook
GW231123 is a puzzling GW detection, with BHs that are among both the most massive and the most spinning detected so far. The primary BH, in particular, can reach $m_{1}\sim 160{\rm M}_{\odot}$, and is thus a promising candidate for a BH “beyond the gap,” with a progenitor subject to the photodisintegration instability.
We performed the first investigation of stars above the upper edge of the BHMG that includes both the effect of stellar rotation and variations in the critical $^{12}\mathrm{C}(\alpha,\gamma){}^{16}\mathrm{O}$ nuclear-reaction rate. The main conclusion of this paper is that stellar evolution alone can explain the primary mass and spin of GW231123, suggesting that this may be the first detection of a BH originating from a photodisintegrated star.
Previous studies have shown that much of the information on population features often comes from a few highly informative events [46, 47, 48, 49, 50]. As was the case for GW190521 when analyzing data from the first three LIGO/Virgo/KAGRA observing runs [51], we anticipate that the exceptionally large masses and spins of GW231123 will play a major role in upcoming GW population constraints with O4 data. Accordingly, our conclusions that the large masses of GW231123 can be produced in combination with large spins through conventional stellar evolution provides a foundation for population‐level studies. We anticipate that it will be important for future GW population analyses to allow the stellar component to accommodate systems like GW231123, rather than imposing a hard cutoff at the BHMG. Fixing such a cutoff a priori could bias inferences toward exotic formation scenarios.
Acknowledgements
We thank Monica Colpi for discussions. D.C. is supported by STFC Grant No. ST/T001011/1. J.S. is supported by NSF Grant No. 2207880. D.G. is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), MSCA Fellowship No. 101064542–StochRewind, MSCA Fellowship No. 101149270–ProtoBH, MUR Young Researchers Grant No. SOE2024-0000125, and the ICSC National Research Centre funded by NextGenerationEU. Our simulations were run on the University of Hawai i’s high-performance supercomputer KOA. The technical support and advanced computing resources from University of Hawai i Information Technology Services – Cyberinfrastructure, funded in part by the NSF MRI award No. 1920304, are gratefully acknowledged.